# Linearization of ergodic McKean SDEs and applications

### v1

@@ Abstract
In this article, we consider McKean stochastic differential equations, as well as their corresponding McKean–Vlasov partial differential equations, which admit a unique stationary state, and we study the linearized Itô diffusion process that is obtained by replacing the law of the process in the convolution term with the unique invariant measure. We show that the law of this linearized process converges to the law of the nonlinear process exponentially fast in time, both in relative entropy and in Wasserstein distance. We study the problem in both the whole space and the torus. We then show how we can employ the resulting linear (in the sense of McKean) Markov process to analyze properties of the original nonlinear and nonlocal dynamics that depend on their long-time behavior. In particular, we propose a linearized maximum likelihood estimator for the nonlinear process which is asymptotically unbiased, and we study the joint diffusive-mean field limit of the underlying interacting particle system.
@@
Key words. Diffusive limit, interacting particle systems, invariant measure, logarithmic Sobolev inequality, maximum likelihood estimation, McKean SDE, McKean–Vlasov PDE, relative entropy.
@@ 1 Introduction
In addition, multiscale problems for interacting particle systems and McKean SDEs have been studied in recent years [32, 22, 4]. In particular, in [22] the joint diffusive-mean field limit for systems of weakly interacting reversible diffusions with periodic confining and interaction potentials was studied. It was shown that the diffusive and mean field limits commute when the mean field system restricted on the torus has a unique invariant measure. On the other hand, it was rigorously established that the two limits do not commute when the dynamics on the torus exhibits phase transitions. Furthermore, the covariance matrix of the limiting Brownian motion can be calculated using the standard homogenization (Kipnis–Varadhan) formula [49, Section 13.6.1], for a linear elliptic cell problem, obtained by calculating the convolution in the generator of the McKean SDE with respect to the Gibbs measure of the dynamics in the torus. In the present work, we show that the the diffusive limit of the interacting particle system and the formula for the diffusion coefficient can be obtained from the linearization procedure.
@@
for some appropriate constant $\lambda>0$. The details can be found in Section 2. The fact that the laws $\mu_{t}$ and $\nu_{t}$ are exponentially close follows from this estimate.
@@
  • We prove that, under assumptions on the drift and confining potential that ensure uniform propagation of chaos, the McKean SDE on $\mathbb{R}^{d}$ is exponentially close in time to the linearized process, both in relative entropy and in the $2$-Wasserstein distance.
@@
The remainder of the paper is organized as follows. In Section 2 we show the convergence analysis for the law of the linearized McKean SDE, which we perform both in the whole space in Section 2.2 and on the torus in Section 2.3. Then, in Section 3 we present two applications of the linearized process: maximum likelihood estimation in Section 3.1 and computation of the asymptotic variance in a central limit theorem in Section 3.2. Finally, in Section 4 we draw our conclusions and outline future developments of the linearization procedure.
@@ 2.1 Preliminaries
The situation is more complicated when we do not have uniqueness of the invariant measure. However, Assumption 2.1 is not a strong limitation of the methodology introduced in this work. In fact, in [44] it is shown that when multiple invariant measures can exist, it is still possible to prove exponentially fast convergence to equilibrium locally, i.e., as long as the initial distribution is sufficiently close to the stationary state and we are away from the critical temperature, where we only get algebraic convergence. Therefore, we believe that the results presented here can be extended to the case of multiple invariant measures by selecting the “correct” density for the linearization procedure depending on basins of attraction of the different invariant measures. In fact, this idea was already employed in the numerical experiments on inference for the Desai–Zwanzig model in [51].
@@ 2.2 Analysis on $\mathbb{R}^{d}$
Under Assumption 2.3, there exists $M>0$ such that for all $t\geq 0$
$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left\lvert\nabla W(x-z)\right\rvert^{2}(f_{s}(z)+f_{\infty}(z))f_{s}(x)\,\mathrm{d}z\,\mathrm{d}x\leq M.$ (2.9)
@@
Let $W$ satisfy Assumption 2.3(iii), then, for all probability density functions $f$ of measures in $\mathcal{P}(\mathbb{R}^{d})$, the function $W*f$ is uniformly convex with constant $\gamma\geq 0$, i.e., it holds for all $x,y\in\mathbb{R}^{d}$
@@
which due to Assumption 2.3(iii) implies
@@ 2.2.1 Convergence of the linearized process
In the classical work [41] as well in the interesting recent paper [43], it is shown that, provided that the initial condition satisfies an LSI, and under our convexity assumptions on the confining and interaction potentials given in Assumption 2.3, then so does the solution of the McKean–Vlasov PDE (1.3), with an LSI constant that can be explicitly computed. In the following lemma we show that the same conclusion holds for the linearized McKean–Vlasov PDE, with the same LSI constant.
@@
Let Assumption 2.3 be satisfied and assume that $\mu_{0}$ and $\nu_{0}$ satisfy an LSI with constant $\lambda_{0}>0$. Then $\mu_{t}$ and $\nu_{t}$ satisfy an LSI with constant
@@
The desired result follows directly from [43, Proposition 1.1] and Lemma 2.5, and the uniform in time constant is obtained by noticing that $\lambda_{t}$ is a convex combination of $\lambda_{0}$ and $1/(2\beta(\alpha+\gamma))$. ∎
@@
Let the assumptions of Lemma 2.6 be satisfied. Then, it holds
@@
which due to Lemma 2.6 implies
@@
which, by [39, Lemma A.1], implies
@@
Let the assumptions of Theorem 2.7 be satisfied. Then, it holds
@@
where $\mathcal{W}_{2}$ denotes the 2-Wasserstein distance and $\rho_{\Lambda}(t)$ is defined in Theorem 2.7.
@@
The desired result follows from Theorem 2.7 using [48, Theorem 1]. ∎
@@
Let Assumption 2.3 be satisfied. Then, it holds
@@
where Lemma 2.5 gives
@@
If the LSI constant $\lambda_{0}$ of the initial distributions $\mu_{0}$ and $\nu_{0}$ is sufficiently small, namely $\lambda_{0}<4/(\alpha\beta)$, then the rate of convergence in Theorems 2.7 and 2.8 is determined only by the convexity parameter $\alpha$ of the confining potential. This is in agreement with the rate of convergence in $L^{2}$ predicted by Theorem 2.9.
@@ 2.3 Analysis on $\mathbb{T}^{d}$
In this section, the domain $\mathbb{D}$ is the multidimensional torus $\mathbb{T}^{d}$ of size length $1$, and we work in the framework below.
@@
Notice that Assumption 2.11(iii) guarantees that Assumption 2.1 is satisfied. In fact, all invariant probability density functions $f_{\infty}$ are also minimizers of the free energy
@@
which is strictly convex under Assumption 2.11(iii). This is shown in [14, Proposition 2.8] without the confining potential, which, however, does not affect the convexity of $\mathcal{E}$ since its corresponding term is linear; see also [3, Section V.B.]. Moreover, the invariant measure $\mu_{\infty}$ satisfies an LSI, as shown in the following result.
@@
Under Assumption 2.11, the invariant measure $\mathrm{d}\mu_{\infty}(x)=f_{\infty}(x)\,\mathrm{d}x$ satisfies an LSI with constant $\Gamma/(2\pi^{2})$, where
@@
We follow [33, Lemma 2.4]. From the integral equation (1.4) we deduce that for all $x\in\mathbb{T}^{d}$
@@
Then, the desired result follows from the Holley–-Stroock perturbation lemma (see, e.g., [33, Lemma 2.3]) and the fact that the uniform distribution on the torus satisfies an LSI (see, e.g., [33, Lemma 2.4]). We remark that the discrepancy with [33, Lemma 2.4] is due to the factor $4$ in the LSI (2.6). ∎
@@ 2.3.1 Convergence to equilibrium for the McKean–Vlasov PDE
Before presenting the convergence results for the linearized process, we need the exponentially fast convergence of the measure $\mu_{t}$ to the invariant measure $\mu_{\infty}$. This problem has already been studied when there is no confining potential in, e.g., [17, 14]. In the absence of confining potential, the analysis is simplified because the unique, under Assumption 2.11(iii), invariant measure is the Lebesgue measure. However, here we extend the results to the case of a nonzero confining potential. Given two probabilities $\mu,\nu\in\mathcal{P}(\mathbb{T}^{d})$ with densities $f,g$, we denote their total variation distance by $\mathrm{TV}(\mu,\nu)$. Moreover, we recall that it holds the following equivalence with the $L^{1}(\mathbb{T}^{d})$ distance between the densities
@@
Let Assumption 2.11 be satisfied. Then, it holds
@@
$\left\lvert I_{t}\right\rvert$ $\leq\left(\left\|\nabla V\right\|_{L^{\infty}(\mathbb{T}^{d})}+\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}\right)\left\|f_{t}-f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}\left\|\nabla(f_{t}-f_{\infty})\right\|_{L^{1}(\mathbb{T}^{d})}$ (2.43)
@@
Let Assumption 2.11 be satisfied. Then, it holds
@@
We now need to bound the terms in the right-hand side. First, by Hölder’s and Young’s convolution inequalities, we have
@@
Finally, Lemma 2.12 yields
@@
In Propositions 2.14 and 2.15, the exponents $\zeta$ and $\eta$ are always strictly positive in the high temperature regime, i.e., when $\beta$ is sufficiently large. Moreover, from the proof of [13, Proposition 3.1] it follows that we can replace $\left\|\Delta W\right\|_{L^{\infty}(\mathbb{T}^{d})}$ in (2.49) with $\left\|\Delta W_{u}\right\|_{L^{\infty}(\mathbb{T}^{d})}$, where $W_{u}$ denotes the non-$H$-stable part of $W$ and is defined in [13, Definition 2.1]. However, the same argument cannot be applied to $\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}$.
The corollary below is a straightforward consequence of Propositions 2.14 and 2.15, and quantifies the rate of convergence in $L^{1}(\mathbb{T}^{d})$.
@@
Let Assumption 2.11 be satisfied. Then, it holds for $i=1,2$
@@
The desired result follows applying Propositions 2.14 and 2.15, the CKP inequality (2.57), and bounding the $L^{1}(\mathbb{T}^{d})$ norm with the $L^{2}(\mathbb{T}^{d})$ one. ∎
@@ 2.3.2 Convergence of the linearized process
Let Assumption 2.11 be satisfied, and assume that there exists a constant $\kappa>1$ such that for all $x\in\mathbb{T}^{d}$
@@
where $\Gamma$ is defined in (2.36). Moreover, if $a_{i}>0$ and $C_{i}/a_{i}<\mathbb{W}(1)$ for either $i=1$ or $i=2$, with $a_{i}$ and $C_{i}$ given in Corollary 2.17 and equation (2.77) below, and where $\mathbb{W}$ denotes the Lambert function, then $\mu_{t}$ satisfies a time-uniform LSI with constant
@@
$+\beta\int_{0}^{t}\operatorname{\mathbb{E}}\left[(\nabla V(Z_{s}^{\mu})+\nabla W*f_{\infty}(Z_{s}^{\mu}))\cdot\nabla W*((\psi_{T-s}-1)f_{\infty})(Z_{s}^{\mu})\psi_{T-s}(Z_{s}^{\mu})\,\mathrm{d}s\right].$
Using Corollary 2.17 we have
@@
Therefore, the desired results follow from the bounds (2.71) and (2.82) by applying the Holley–Stroock perturbation lemma (see, e.g., [33, Lemma 2.3]) and Lemma 2.12. ∎
Using the LSI, we can finally prove that the measures $\mu_{t}$ and $\nu_{t}$ are exponentially close in time, in relative entropy. The argument is similar to the proof of Theorem 2.7.
@@
Let the assumptions of Lemma 2.18 be satisfied. Then, it holds for $i=1,2$
@@
where $a_{i}$ and $c_{i}$ are defined in Corollary 2.17.
@@
which, due to Lemma 2.18, implies
@@
Then, Corollary 2.17 gives
@@
Applying [39, Lemma A.1], we have
@@
Let the assumptions of Theorem 2.19 be satisfied. Then, it holds
@@
where $\mathcal{W}_{2}$ denotes the 2-Wasserstein distance and $\sigma_{\Xi}(t)$ is defined in Theorem 2.19.
@@
The desired result follows from Theorem 2.19 using [48, Theorem 1]. ∎

### v2

@@ Abstract
In this article, we consider McKean stochastic differential equations, as well as their corresponding McKean–Vlasov partial differential equations, which admit a unique stationary state, and we study the linearized Itô diffusion process that is obtained by replacing the law of the process in the convolution term with the unique invariant measure. We show that the law of the nonlinear McKean process converges to the law of this linearized process exponentially fast in time, both in relative entropy and in Wasserstein distance. We study the problem in both the whole space and the torus. We then show how we can employ the resulting linear (in the sense of McKean) Markov process to analyze properties of the original nonlinear and nonlocal dynamics that depend on their long-time behavior. In particular, we propose a linearized maximum likelihood estimator for the nonlinear process which is asymptotically unbiased, and we study the joint diffusive-mean field limit of the underlying interacting particle system.
@@
Key words. Diffusive limit, interacting particle systems, invariant measure, logarithmic Sobolev inequality, maximum likelihood estimation, McKean–Vlasov PDE.
@@ 1 Introduction
In addition, multiscale problems for interacting particle systems and McKean SDEs have been studied in recent years [32, 22, 4]. In particular, in [22] the joint diffusive-mean field limit for systems of weakly interacting reversible diffusions with periodic confining and interaction potentials was studied. It was shown that the diffusive and mean field limits commute when the mean field system restricted on the torus has a unique invariant measure. On the other hand, it was rigorously established that the two limits do not commute when the dynamics on the torus exhibits phase transitions. Furthermore, the covariance matrix of the limiting Brownian motion can be calculated using the standard homogenization (Kipnis–Varadhan) formula [49, Section 13.6.1], for a linear elliptic cell problem, obtained by calculating the convolution in the generator of the McKean SDE with respect to the Gibbs measure of the dynamics in the torus. In the present work, we show that the diffusive limit of the interacting particle system and the formula for the diffusion coefficient can be obtained from the linearization procedure.
@@
for some appropriate constant $\lambda>0$. The details can be found in Section˜2. The fact that the laws $\mu_{t}$ and $\nu_{t}$ are exponentially close follows from this estimate.
@@
  • We prove that, under assumptions on the confining and interaction potentials that ensure uniform propagation of chaos, the McKean SDE on $\mathbb{R}^{d}$ is exponentially close in time to the linearized process, both in relative entropy and in the 222-Wasserstein distance.
@@
The remainder of the paper is organized as follows. In Section˜2 we show the convergence analysis for the law of the linearized McKean SDE, which we perform both in the whole space in Section˜2.2 and on the torus in Section˜2.3. Then, in Section˜3 we present two applications of the linearized process: maximum likelihood estimation in Section˜3.1 and computation of the asymptotic variance in a central limit theorem in Section˜3.2. Finally, in Section˜4 we draw our conclusions and outline future developments of the linearization procedure.
@@ 2.1 Preliminaries
The situation is more complicated when we do not have uniqueness of the invariant measure. However, Assumption˜2.1 is not a strong limitation of the methodology introduced in this work. In fact, in [44] it is shown that when multiple invariant measures can exist, it is still possible to prove exponentially fast convergence to equilibrium locally, i.e., as long as the initial distribution is sufficiently close to the stationary state and we are away from the critical temperature, where we only get algebraic convergence. Therefore, we believe that the results presented here can be extended to the case of multiple invariant measures by selecting the “correct” density for the linearization procedure depending on basins of attraction of the different invariant measures. In fact, this idea was already employed in the numerical experiments on inference for the Desai–Zwanzig model in [51].
@@ 2.2 Analysis on $\mathbb{R}^{d}$
Under Assumption˜2.3, there exists $M>0$ such that for all $t\geq 0$
$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\left\lvert\nabla W(x-z)\right\rvert^{2}(f_{t}(z)+f_{\infty}(z))f_{t}(x)\,\mathrm{d}z\,\mathrm{d}x\leq M.$ (2.9)
@@
Let $W$ satisfy Assumption˜2.3(iii), then, for all probability density functions $f$ of measures in $\mathcal{P}(\mathbb{R}^{d})$, the function $W*f$ is uniformly convex with constant $\gamma\geq 0$, i.e., it holds for all $x,y\in\mathbb{R}^{d}$
@@
which due to Assumption˜2.3(iii) implies
@@ 2.2.1 Convergence of the linearized process
In the classical work [41] as well in the interesting recent paper [43], it is shown that, provided that the initial condition satisfies an LSI, and under our convexity assumptions on the confining and interaction potentials given in Assumption˜2.3, then so does the solution of the McKean–Vlasov PDE (1.3), with an LSI constant that can be explicitly computed. In the following lemma we show that the same conclusion holds for the linearized McKean–Vlasov PDE, with the same LSI constant.
@@
Let Assumption˜2.3 be satisfied and assume that $\mu_{0}$ and $\nu_{0}$ satisfy an LSI with constant $\lambda_{0}>0$. Then $\mu_{t}$ and $\nu_{t}$ satisfy an LSI with constant
@@
The desired result follows directly from [43, Proposition 1.1] and Lemma˜2.5, and the uniform in time constant is obtained by noticing that $\lambda_{t}$ is a convex combination of $\lambda_{0}$ and $1/(2\beta(\alpha+\gamma))$. ∎
@@
Let the assumptions of Lemma˜2.6 be satisfied. Then, it holds
@@
which due to Lemma˜2.6 implies
@@
which, by the Grönwall’s inequality in [39, Lemma A.1], implies
@@
Let the assumptions of Theorem˜2.7 be satisfied. Then, it holds
@@
where $\mathcal{W}_{2}$ denotes the 2-Wasserstein distance and $\rho_{\Lambda}(t)$ is defined in Theorem˜2.7.
@@
The desired result follows from Theorem˜2.7 using [48, Theorem 1]. ∎
@@
Let Assumption˜2.3 be satisfied. Then, it holds
@@
where Lemma˜2.5 gives
@@
If the LSI constant $\lambda_{0}$ of the initial distributions $\mu_{0}$ and $\nu_{0}$ is sufficiently small, namely $\lambda_{0}<4/(\alpha\beta)$, then the rate of convergence in Theorems˜2.7 and 2.8 is determined only by the convexity parameter $\alpha$ of the confining potential. This is in agreement with the rate of convergence in $L^{2}$ predicted by Theorem˜2.9.
@@ 2.3 Analysis on $\mathbb{T}^{d}$
In this section, the domain $\mathbb{D}$ is the multidimensional torus $\mathbb{T}^{d}$ of size length 111, and we work in the framework below.
@@
Notice that Assumption˜2.11(iii) guarantees that Assumption˜2.1 is satisfied. In fact, all invariant probability density functions $f_{\infty}$ are also minimizers of the free energy
@@
which is strictly convex under Assumption˜2.11(iii). This is shown in [14, Proposition 2.8] without the confining potential, which, however, does not affect the convexity of $\mathcal{E}$ since its corresponding term is linear; see also [3, Section V.B.]. Moreover, the invariant measure $\mu_{\infty}$ satisfies an LSI, as shown in the following result.
@@
Under Assumption˜2.11, the invariant measure $\mathrm{d}\mu_{\infty}(x)=f_{\infty}(x)\,\mathrm{d}x$ satisfies an LSI with constant $\Gamma/(2\pi^{2})$, where
@@
We follow [33, Lemma 2.4]. From the integral equation (1.4) and the Young’s convolution inequality we deduce that for all $x\in\mathbb{T}^{d}$
$f_{\infty}(x)=\frac{e^{-\beta(V(x)+(W*f_{\infty})(x))}}{\int_{\mathbb{T}^{d}}e^{-\beta(V(z)+(W*f_{\infty})(z))}\,\mathrm{d}z}\leq\frac{e^{\beta\left(\left\|V\right\|_{L^{\infty}(\mathbb{T}^{d})}+\left\|W\right\|_{L^{\infty}(\mathbb{T}^{d})}\left\|f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}\right)}}{\int_{\mathbb{T}^{d}}e^{-\beta\left(\left\|V\right\|_{L^{\infty}(\mathbb{T}^{d})}+\left\|W\right\|_{L^{\infty}(\mathbb{T}^{d})}\left\|f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}\right)}\,\mathrm{d}z}.$ (2.37)
Proceeding analogously for the lower bound, and noting that $\left\|f_{\infty}\right\|_{L^{1}(\mathbb{T}^{d})}=1$ since $f_{\infty}$ is a probability density function and that $\mathbb{T}^{d}$ has Lebesgue measure 1, we obtain
@@
Then, the desired result follows from the Holley–-Stroock perturbation lemma (see, e.g., [33, Lemma 2.3]) and the fact that the uniform distribution on the torus satisfies an LSI (see, e.g., [33, Lemma 2.4]). We remark that the discrepancy with [33, Lemma 2.4] is due to the factor 444 in the LSI (2.6). ∎
@@ 2.3.1 Convergence to equilibrium for the McKean–Vlasov PDE
Before presenting the convergence results for the linearized process, we need the exponentially fast convergence of the measure $\mu_{t}$ to the invariant measure $\mu_{\infty}$. This problem has already been studied when there is no confining potential in, e.g., [17, 14]. In the absence of confining potential, the analysis is simplified because the unique, under Assumption˜2.11(iii), invariant measure is the Lebesgue measure. However, here we extend the results to the case of a nonzero confining potential. Given two probabilities $\mu,\nu\in\mathcal{P}(\mathbb{T}^{d})$ with densities $f,g$, we denote their total variation distance by $\mathrm{TV}(\mu,\nu)$. Moreover, we recall the following equivalence with the $L^{1}(\mathbb{T}^{d})$ distance between the densities (see, e.g., [61, Lemma 2.1])
@@
Let Assumption˜2.11 be satisfied. Then, it holds
@@
$\left\lvert I_{t}\right\rvert$ $\leq\left(\left\|\nabla V\right\|_{L^{\infty}(\mathbb{T}^{d})}+\left\|\nabla W\right\|_{L^{\infty}(\mathbb{T}^{d})}\right)\left\|f_{t}-f_{\infty}\right\|_{L^{2}(\mathbb{T}^{d})}\left\|\nabla(f_{t}-f_{\infty})\right\|_{L^{2}(\mathbb{T}^{d})}$ (2.44)
@@
Let Assumption˜2.11 be satisfied. Then, it holds
@@
We now need to bound the terms in the right-hand side. First, by Hölder’s and Young’s convolution inequalities and since $\nabla\log f_{\infty}=-\beta(\nabla V+\nabla W*f_{\infty})$, we have
@@
Finally, Lemma˜2.12 yields
@@
[…]
The corollary below is a straightforward consequence of Propositions˜2.14 and 2.15, and quantifies the rate of convergence in $L^{1}(\mathbb{T}^{d})$.
@@
Let Assumption˜2.11 be satisfied. Then, it holds for $i=1,2$
@@
The desired result follows applying Propositions˜2.14 and 2.15, the CKP inequality (2.58), and bounding the $L^{1}(\mathbb{T}^{d})$ norm with the $L^{2}(\mathbb{T}^{d})$ one. ∎
@@ 2.3.2 Convergence of the linearized process
Let Assumption˜2.11 be satisfied, and assume that there exists a constant $\kappa>1$ such that for all $x\in\mathbb{T}^{d}$
@@
where $\Gamma$ is defined in (2.36). Moreover, if $a_{i}>0$ and $C_{i}/a_{i}<\mathbb{W}(1)$ for either $i=1$ or $i=2$, with $a_{i}$ and $C_{i}$ given in Corollary˜2.17 and equation (2.78) below, and where $\mathbb{W}$ denotes the Lambert function, then $\mu_{t}$ satisfies a time-uniform LSI with constant
@@
$\hskip-14.22636pt+\beta\int_{0}^{t}\operatorname{\mathbb{E}}\left[(\nabla V(Z_{s}^{\mu})+\nabla W*f_{\infty}(Z_{s}^{\mu}))\cdot\nabla W*((\psi_{T-s}-1)f_{\infty})(Z_{s}^{\mu})\psi_{T-s}(Z_{s}^{\mu})\,\mathrm{d}s\right].$
Using Corollary˜2.17 we have
@@
Therefore, the desired results follow from the bounds (2.72) and (2.83) by applying the Holley–Stroock perturbation lemma (see, e.g., [33, Lemma 2.3]) and Lemma˜2.12. ∎
Using the LSI, we can finally prove that the measures $\mu_{t}$ and $\nu_{t}$ are exponentially close in time, in relative entropy. The argument is similar to the proof of Theorem˜2.7.
@@
Let the assumptions of Lemma˜2.18 be satisfied. Then, it holds for $i=1,2$
@@
where $a_{i}$ and $c_{i}$ are defined in Corollary˜2.17.
@@
which, due to Lemma˜2.18, implies
@@
Then, Corollary˜2.17 gives
@@
Applying the Grönwall’s inequality in [39, Lemma A.1], we have
@@
Let the assumptions of Theorem˜2.19 be satisfied. Then, it holds
@@
where $\mathcal{W}_{2}$ denotes the 2-Wasserstein distance and $\sigma_{\Xi}(t)$ is defined in Theorem˜2.19.
@@
The desired result follows from Theorem˜2.19 using [48, Theorem 1]. ∎